# VisionLLM v2: An End-to-End Generalist Multimodal Large Language Model for Hundreds of Vision-Language Tasks

### v2

@@ 2.1 Multimodal Large Language Model
Extension of MLLMs’ Text Output. To extend MLLMs to downstream tasks, models like Kosmo-2 [137], Shikra [25], VisionLLM [176], Ferret [195, 204], and All-Seeing V2 [174] achieve this using specially-designed tokens or encoding coordinates as text tokens. Despite these advancements, using LLMs solely as visual decoders falls short of resolving the fine-grained visual context needed for precise detection and segmentation.
@@ 3.2 Super Link Technique
Visual Perception covers a wide range of visual tasks, such as open-ended/closed-set object detection, instance segmentation, pose estimation, etc. VisionLLM v2 supports using both text and visual prompts to define these tasks. We list an example in the following figure. <image> and <region> are the placeholders that will be replaced by image and region embeddings before being fed into the LLM. Here, we take Example 1 of interactive segmentation for clarification. The user prompts the model to segment specific regions within a question. MLLM sequentially lists the region names followed by a routing token [SEG] in the response. Remember that the proposed method would automatically append the super-link queries after the routing token. In that way, we can obtain the per-region representations by extracting the output hidden states of MLLM from corresponding super-link queries and pooling them into one embedding. These embeddings are fed into a segmentation decoder as the conditional feature, requiring only a single forward to produce segmentation results for all regions.
Example: Text Prompt + Visual Prompt for Interactive Segmentation. USER: <image> Could you please segment all the corresponding objects according to the visual prompts as region1 <region>, region2 <region>? ASSISTANT: Sure, these objects are region1 [SEG], region2 [SEG].
Visual Generation is also a wide topic covering a number of different tasks, such as generation, editing, variation, personalization, etc. In VisionLLM v2, we focus on two fundamental tasks, i.e., text-to-image generation and instruction-based image editing. We use Stable Diffusion v1.5 (SD) as our tool in the text-to-image generation task. We abandon its text encoder and use the output hidden states of the MLLM as the image generation condition for SD. Image editing task [81] can also be accomplished in the same paradigm by using both image and text prompts as inputs.
@@ 3.3 Training Strategy
Stage-2: Multi-capacity Fine-tuning. At this stage, we integrate task-specific decoders into the model and perform multi-task joint training. In addition to the instruction data utilized in stage-1, we incorporate extensive visual datasets such as COCO [102], ADE20K [214] for their specific tasks. We construct a series of instruction templates for these visual datasets to perform instruction tuning, ensuring that the LLM can accurately invoke the downstream decoders. During this stage, the region encoder and all decoders undergo training, and the LLM is finetuned using LoRA [65].
@@ 4.2 Mutimodal Benchmarks
VisionLLM v2 CLIP-L Vicuna-7B 80.8   64.8 51.3   94.7 63.6 88.0 1486.2 77.1 / 66.4 65.3 / 71.5 / 41.9
Table 1: Comparison with SoTA models on multimodal dialogue benchmarks. The academic-oriented datasets include: VQAv2 test-dev [55], GQA test-balanced [70], VizWiz test-dev [60], ScienceQA test [151] and TextVQA val [156]. The instruction-following datasets include: POPE [99], MME [47], MMBench-EN/CN [112], SEED-Bench (all/image/video) [53] and MM-Vet [199]. The training annotations of the dataset are observed during training.
@@ 4.3 Visual Perception Tasks
VisionLLM v2 Generalist Swin-T 56.3 74.3 61.6 47.8 71.8 52.0 92.6 44.2 98.2
@@
VisionLLM v2 Generalist Swin-T 74.2 79.8 77.8 72.6 83.1 93.2 95.5 99.4 99.9
@@
Object Detection and Instance Segmentation. In Table 3, we compare the results of VisionLLM v2 with state-of-the-art methods on two fundamental vision tasks, i.e., object detection, and instance segmentation. As can be seen, using the lightweight backbone Swin-T, our generalist model achieves the performance of 56.3 APb and 47.8 APm on COCO. The results significantly outperform the previous methods using ResNet50 [62] backbone and are comparable with the specialist model Grounding-DINO-T [110]. Moreover, we also validate our model on the crowded pedestrian detection dataset, i.e., CrowdHuman. VisionLLM v2 surpasses the previous best generalist model Hulk [179] by 0.4 points on AP50.
@@ A.1 More Experimental Results
Visual Grounding. Visual grounding is a crucial vision task that associates the language description with the specific object within an image. Using the box or mask as the output format, visual grounding can be further categorized into referring expression comprehension (REC) and referring expression segmentation (RES) tasks. We comprehensively list the comparison results for the two tasks in Table A2 and Table A3, respectively. From Table A2, it is found that VisionLLM v2 achieves the best performance on RefCOCO [198] among MLLMs. We also showcase that VisionLLM v2 exhibits remarkable pixel-level segmentation capacities in Table A3. Without further fine-tuning, our model demonstrates the best gIoU result of 56.9 on the challenging ReasonSeg dataset [87].
@@
method type val testA testB val testA testB val test gIoU cIoU
X-Decoder (L) [220] - - - - - - 64.6 - - -
SEEM (L) [221] - - - - - - 65.6 - - -
UNINEXT (R50) [189] 77.9 79.7 75.8 66.2 71.2 59.0 70.0 70.5 - -
UNINEXT (H) [189] 82.2 83.4 81.3 72.5 76.4 66.2 74.7 76.4 - -
GLEE-Pro [184] VGM 80.0 - - 69.6 - - 72.9 - - -
LISA-7B [87] 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5 44.4 46.0
LISA-7B [87] 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6 52.9 54.0
PixelLM [148] 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5 - -
PixelLLM [187] 76.9 78.5 74.4 69.2 72.1 64.5 70.7 72.4 - -
AnyRef [61] 76.9 79.9 74.2 70.3 73.5 61.8 70.0 70.7 - -
GROUNDINGHOG [209] 78.5 79.9 75.7 70.5 75.0 64.9 74.1 74.6 56.2 -
GLaMM [145] 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9 - -
VisionLLM v2 MLLM 79.2 82.3 77.0 68.9 75.8 61.8 73.3 74.8 56.9 48.3
Table A3: Comparison of referring expression segmentation performance. gIoU and cIoU denote the general IoU and cumulative IoU, respectively. The results for RefCOCO/+/g [198, 123] are reported based on cIoU. VGM and MLLM represent vision generalist models and multimodal large language models, respectively. The model is finetuned on the dataset.
@@
Camouflaged Object Detection. The performance comparisons for camouflaged object detection (COD) are presented in Table A7. It is observed that VisionLLM v2 exhibits competitive performance with state-of-the-art expert models that undergo longer training schedule, e.g., 150 epochs, on specific datasets.
@@ A.3 Zero-shot Evaluation
Table A9: Ablation on the comparison between super-link queries and token embeddings. We evaluate the results on the four crucial visual perception tasks: instance segmentation (inst seg.), visual grounding (ground.), pose estimation (pose) and interactive segmentation (interact seg.). Our default setting is marked in gray.
@@ Appendix D Training Details
Settings of Stage-2. In stage-2, we add the task-specific decoders and perform the multi-capacity fine-tuning. The LLM is equipped with LoRA [65] for saving computational memory, using a LoRA rank of 32 and an alpha of 64. LLM and region encoder are trained with the peak learning rate of 1e-5, while the decoders are trained with the peak learning rate of 1e-4. The model is trained on 128 A100 GPUs with a batch size of 2 per GPU.
@@ 7. Can you paint a vivid picture of the scenery within <regions> captured in the image?
1. Can you provide me with a brief description of <regions> in the picture? 2. I’m curious about the region represented by <regions> in the picture. Could you describe it in short? 3. What can you tell me about <regions> in the image? 4. I’d like to know more about the area in the photo labeled <regions>. Can you give me a brief description? 5. Could you describe <regions> in the picture in short? 6. What content can you give me about <regions> in the photo? 7. Please provide me with a short description of <regions> in the image. 8. Can you give me a brief account of the region labeled as <regions> in the picture? 9. I’m interested in learning more about <regions> in the photo. Can you describe it in short? 10. What is the region outlined by <regions> in the picture like? Could you give me a brief description? 11. Can you provide me with a brief description of <regions> in the picture, please? 12. I’m curious about the region represented by <regions> in the picture. Could you describe it in short, please? 13. What can you tell me about <regions> in the image, exactly? 14. I’d like to know more about <regions>. Can you give me a brief description? 15. Could you describe the region shown as <regions> in the picture in short, please? 16. What content can you give me about <regions> in the photo, please? 17. Please provide me with a short description of <regions> in the image, please. 18. Can you give me a brief account of the region labeled as <regions> in the picture, please? 19. I’m interested in learning more about <regions> in the photo. Can you describe it in short, please? 20. What is <regions> in the picture like, please? Could you give me a brief description?
@@ 30. Can you provide a brief explanation of the specific locations marked by <regions> in this photo?
1. Whis is the object category of <regions>? Answer the question with a single word or phrase. 2. Could you tell me what is the object in <regions>? Answer the question with a single word or phrase. 3. What category best describes the area represented by <regions>? Answer the question with a single word or phrase. 4. Can you specify the type of object inside the region labeled by <regions>? Answer the question with a single word or phrase. 5. How would you label the area indicated by <regions> in the image? Answer the question with a single word or phrase. 6. Give a category label to the region outlined by <regions>. Answer the question with a single word or phrase. 7. Please identify the category of the object inside the <regions>. Answer the question with a single word or phrase. 8. Examine and determine the primary subject located within <regions>. Answer the question with a single word or phrase. 9. I need your help to assign an object category to the <regions>, please. Answer the question with a single word or phrase. 10. Evaluate the content of the region shown as <regions> and provide its category. Answer the question with a single word or phrase.

### v3

@@ 2.1 Multimodal Large Language Model
Extension of MLLMs’ Text Output. To extend MLLMs to downstream tasks, models like Kosmo-2 [140], Shikra [27], VisionLLM [182], Ferret [201, 212], and All-Seeing V2 [180] achieve this using specially-designed tokens or encoding coordinates as text tokens. Despite these advancements, using LLMs solely as visual decoders falls short of resolving the fine-grained visual context needed for precise detection and segmentation. The other line of works focus on broadening the modality scope. AnyGPT [210] builds a multimodal text-centric dataset for any-to-any multimodal generation (text, image, speech, music) with sequence modeling. Chameleon [168] uses fully token-based representations for both texts and images, capable of understanding and generating interleaved image-text sequences. CM3leon [5, 205] are autoregressive models for text-to-image and image-to-text tasks. All these works could unify image understanding and generation in one network. Our model can support more vision and vision-language tasks.
@@ 3.2 Super Link Technique
Visual Perception covers a wide range of visual tasks, such as open-ended/closed-set object detection, instance segmentation, pose estimation, etc. VisionLLM v2 supports using both text and visual prompts to define these tasks. We list an example in the following figure. <image> and <region> are the placeholders that will be replaced by image and region embeddings before being fed into the LLM. Here, we take Example 1 of interactive segmentation for clarification. The user prompts the model to segment specific regions within a question. MLLM sequentially lists the region names followed by a routing token [SEG] in the response. Remember that the proposed method would automatically append the super-link queries after the routing token. In that way, we can obtain the per-region representations by extracting the output hidden states of MLLM from corresponding super-link queries and pooling them into one embedding. These embeddings are fed into a segmentation decoder as the conditional feature, requiring only a single forward to produce segmentation results for all regions. In the following, we show a template example for interactive segmentation.
Example1: Text Prompt + Visual Prompt for Interactive Segmentation. USER: <image> Could you please segment all the corresponding objects according to the visual prompts as region1 <region>, region2 <region>? ASSISTANT: Sure, these objects are region1 [SEG], region2 [SEG].
Visual Generation is also a wide topic covering a number of different tasks, such as generation, editing, variation, personalization, etc. In VisionLLM v2, we focus on two fundamental tasks, i.e., text-to-image generation and instruction-based image editing. We use Stable Diffusion v1.5 (SD) as our tool in the text-to-image generation task. We abandon its text encoder and use the output hidden states of the MLLM as the image generation condition for SD. Image editing task [82] can also be accomplished in the same paradigm by using both image and text prompts as inputs. In the following, we list a template example for text-to-image generation.
Example 2: Text Prompt for Text-to-Image Generation. ASSISTANT: Of course, here it is [GEN].
Discussion. Some previous works have used the special token or learnable queries independently. InstructBLIP [42], ep-ALM [158], and MAPL [125] use learnable queries (i.e., soft prompts) to connect the modality encoders and LLM. FROMAGe [84] uses a special token for image-text retrieval so as to handle multimodal outputs, where the images are not generated from the network end-to-end. However, these works still remain constrained to text-based outputs. The proposed super link is the seamless integration of the two techniques. Despite the simplicity of our method, it is able to extend MLLMs to handle hundreds of tasks by largely extending the output formats, e.g., box, mask, keypoint and image. Meanwhile, it can address several challenges when scaling up various tasks: (i) precise decoder invocation, (ii) mitigating task conflicts and (iii) efficient message transmission in an end-to-end manner.
@@ 3.3 Training Strategy
Stage-2: Multi-capacity Fine-tuning. At this stage, we integrate task-specific decoders into the model and perform multi-task joint training. In addition to the instruction data utilized in stage-1, we incorporate extensive visual datasets such as COCO [104], ADE20K [222] for their specific tasks. We construct a series of instruction templates for these visual datasets to perform instruction tuning, ensuring that the LLM can accurately invoke the downstream decoders. During this stage, the region encoder and all decoders undergo training, and we only finetune the input and output embeddings of the LLM to maximally preserve its original conversational ability.
@@ 4.2 Mutimodal Benchmarks
VisionLLM v2 CLIP-L Vicuna-7B 80.8   65.1 51.8   94.2 64.7 88.8 1495.6 76.3 / 66.8 65.6 / 71.7 / 42.2
Table 1: Comparison with SoTA models on multimodal dialogue benchmarks. The academic-oriented datasets include: VQAv2 test-dev [57], GQA test-balanced [71], VizWiz test-dev [62], ScienceQA test [154] and TextVQA val [160]. The instruction-following datasets include: POPE [101], MME [49], MMBench-EN/CN [114], SEED-Bench (all/image/video) [55]. The training annotations of the dataset are observed during training.
@@ 4.3 Visual Perception Tasks
VisionLLM v2 Generalist Swin-T 56.7 74.5 62.2 47.8 71.8 52.0 93.1 44.7 98.5
@@
VisionLLM v2 Generalist Swin-T 74.0 79.4 76.8 72.9 81.9 91.1 94.4 99.4 97.8
@@
Object Detection and Instance Segmentation. In Table 3, we compare the results of VisionLLM v2 with state-of-the-art methods on two fundamental vision tasks, i.e., object detection, and instance segmentation. As can be seen, using the lightweight backbone Swin-T, our generalist model achieves the performance of 56.7 APb and 47.8 APm on COCO. The results significantly outperform the previous methods using ResNet50 [64] backbone and are comparable with the specialist model Grounding-DINO-T [112]. Moreover, we also validate our model on the crowded pedestrian detection dataset, i.e., CrowdHuman. VisionLLM v2 surpasses the previous best generalist model Hulk [185] by 0.9 points on AP50.
@@ 4.5 Ablation Study
Train / Test Image VQA Inst. Seg Image Gen.
Image VQA -0.01 -0.11 -0.04
Inst Seg. +0.04 -0.12 + 0.19
Image Gen. +0.03 +0.02 -0.04
Table 6: Ablation on the multi-task influence. The numbers denote the loss change when the model is finetuned on a single task.
TextVQA MME MMB EN/CN COCO COCO-Pose
One-stage 53.2 1284.4 61.9 / 51.4 54.9 / 44.6 74.1
Three-stage 66.2 1507.1 77.8 / 68.5 56.3 / 47.6 74.2
Table 7: Ablation on the one-stage and three-stage training. We evaluate the models on image VQA, instance segmentation and pose estimation.
Multi-Task Influence. As indicated by previous works [225, 206], different tasks with shared parameters may cause conflict with each other. This is mainly due to inconsistent optimization in multi-task learning. To investigate the mutual influence of multi-task joint training in our framework, we start from the same checkpoint and train the model on a single task (image VQA, instance segmentation, or image generation) for 1000 iterations. We record the loss change for all three tasks in Table 7. In the table, a decrease in the loss value indicates beneficial training for the task, while an increase is detrimental. We can observe that training on image VQA is advantageous for all three tasks, which is reasonable as the conversation ability of MLLM is enhanced. Whereas training exclusively on instance segmentation or image generation leads to conflicts with other tasks. This aligns with the findings in Uniperceiver-MoE [225].
One-Stage vs.Three-Stage Training. Some previous generalist models [176, 15] train the model in one stage. Our model encompasses much more tasks and thus introduces a training conflict: the MLLM requires only 1 epoch of training on chat data to prevent overfitting, whereas the decoders need longer training epochs (e.g., Grounding-DINO need 12 epochs of training on visual data) to achieve convergence. One possible solution for one-stage training is to give a higher sample ratio for the visual data. In the following, we conduct the ablation to study the effect of one-stage v.s. three-stage training. We use image-level chat data, COCO, and COCO-Pose for image understanding, instance segmentation, and pose estimation, respectively. For one-stage training, we repeat the COCO and COCO-Pose datasets 12 times. As can be seen from Table 7, the conversation ability of the model is significantly decreased due to extreme data imbalance. And the performance of instance segmentation and pose estimation is also slightly reduced. These results prove the effectiveness of our three-stage training.
@@ 5 Conclusion & Limitation
Acknowledgement
This paper is supported by the National Key R&D Program of China (No.2022ZD0161000), the General Research Fund of Hong Kong (No.17200622, 17209324), and the National Natural Science Foundation of China (No. 62376134, 62372223). Tong Lu and Zhe Chen are supported by the China Mobile Zijin Innovation Insititute (No. NR2310J7M). Zhe Chen is also supported by the Youth PhD Student Research Project under the National Natural Science Foundation (No. 623B2050).
@@ A.1 More Experimental Results
Visual Grounding. Visual grounding is a crucial vision task that associates the language description with the specific object within an image. Using the box or mask as the output format, visual grounding can be further categorized into referring expression comprehension (REC) and referring expression segmentation (RES) tasks. We comprehensively list the comparison results for the two tasks in Table A2 and Table A3, respectively. From Table A2, it is found that VisionLLM v2 achieves the best performance on RefCOCO [204] among MLLMs. We also showcase that VisionLLM v2 exhibits remarkable pixel-level segmentation capacities in Table A3. Without further fine-tuning, our model demonstrates the good gIoU result of 51.0 on the challenging ReasonSeg dataset [89].
@@
method type val testA testB val testA testB val test gIoU
X-Decoder (L) [229] - - - - - - 64.6 - -
SEEM (L) [230] - - - - - - 65.6 - -
UNINEXT (R50) [195] 77.9 79.7 75.8 66.2 71.2 59.0 70.0 70.5 -
UNINEXT (H) [195] 82.2 83.4 81.3 72.5 76.4 66.2 74.7 76.4 -
GLEE-Pro [190] VGM 80.0 - - 69.6 - - 72.9 - -
LISA-7B [89] 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5 44.4
LISA-7B [89] 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6 52.9
PixelLM [151] 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5 -
PixelLLM [193] 76.9 78.5 74.4 69.2 72.1 64.5 70.7 72.4 -
AnyRef [63] 76.9 79.9 74.2 70.3 73.5 61.8 70.0 70.7 -
GROUNDINGHOG [217] 78.5 79.9 75.7 70.5 75.0 64.9 74.1 74.6 56.2
GLaMM [148] 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9 -
VisionLLM v2 MLLM 76.6 79.3 74.3 64.5 69.8 61.5 70.7 71.2 51.0
Table A3: Comparison of referring expression segmentation performance. gIoU denotes the general IoU. The results for RefCOCO/+/g [204, 126] are reported based on cumulative IoU (cIoU). VGM and MLLM represent vision generalist models and multimodal large language models, respectively. The model is finetuned on the dataset.
@@
Camouflaged Object Detection. The performance comparisons for camouflaged object detection (COD) are presented in Table A7. It is observed that VisionLLM v2 exhibits competitive performance with state-of-the-art expert models that undergo longer training schedule, e.g., 150 epochs.
@@ A.3 Zero-shot Evaluation
In-Context Segmentation
Method mIoU
Painter [183] 44.26
SegGPT [184] 54.25
VisionLLM v2 67.51
In-Context Image Captioning
Method METEROR / CIDEr
OpenFlamingo [11] 13.80 / 104.61
VisionLLM v2 18.56 / 152.74
Table A9: Comparison of in-context segmentation and in-context image captioning performance.
In-Context Segmentation & In-Context Image Captioning. To evaluate the in-context learning ability of VisionLLM v2, we compare the results of in-context segmentation and in-context image captioning in Table A9. For in-context segmentation, we construct a benchmark based on the validation set of COCO2017, where the number of in-context examples used during inference ranges from 1 to 5. For in-context image-captioning, we follow the same evaluation protocol as OpenFlamingo [11] and use 4-shot to assess the performance between different methods. The validation set is built upon COCO2017. From the table, VisionLLM v2 exhibits clear performance advantages compared with state-of-the-art methods in both in-context learning settings, which demonstrates the superior in-context capacities of our method.
@@
Table A10: Ablation on the comparison between super-link queries and token embeddings. We evaluate the results on the four crucial visual perception tasks: instance segmentation, visual grounding, pose estimation and interactive segmentation. Our default setting is marked in gray.
@@ Appendix D Training Details
Settings of Stage-2. In stage-2, we add the task-specific decoders and perform the multi-capacity fine-tuning. LLM and region encoder are trained with the peak learning rate of 1e-5, while the decoders are trained with the peak learning rate of 1e-4. The model is trained on 128 A100 GPUs with a batch size of 2 per GPU.
@@
Training Losses. During training, we use the standard cross-entropy loss in stage-1. In stage-2 and stage-3, when integrating the task-specific decoders, we simply sum the losses from the LLM and decoders directly, without reweighting each component. i.e.,
$L_{\text{total}}=L_{\text{llm}}+L_{\text{gdino}}+L_{\text{unipose}}+L_{\text{sd}}+L_{\text{ip2p}}$ (1)
@@ 7. Can you paint a vivid picture of the scenery within <regions> captured in the image?
1. Can you provide me with a brief description of <regions> in the picture?
2. I’m curious about the region represented by <regions> in the picture. Could you describe it in short?
3. What can you tell me about <regions> in the image?
4. I’d like to know more about the area in the photo labeled <regions>. Can you give me a brief description?
5. Could you describe <regions> in the picture in short?
6. What content can you give me about <regions> in the photo?
7. Please provide me with a short description of <regions> in the image.
8. Can you give me a brief account of the region labeled as <regions> in the picture?
9. I’m interested in learning more about <regions> in the photo. Can you describe it in short?
10. What is the region outlined by <regions> in the picture like? Could you give me a brief description?
11. Can you provide me with a brief description of <regions> in the picture, please?
12. I’m curious about the region represented by <regions> in the picture. Could you describe it in short, please?
13. What can you tell me about <regions> in the image, exactly?
14. I’d like to know more about <regions>. Can you give me a brief description?
15. Could you describe the region shown as <regions> in the picture in short, please?
16. What content can you give me about <regions> in the photo, please?
17. Please provide me with a short description of <regions> in the image, please.
18. Can you give me a brief account of the region labeled as <regions> in the picture, please?
19. I’m interested in learning more about <regions> in the photo. Can you describe it in short, please?
20. What is <regions> in the picture like, please? Could you give me a brief description?
@@ 30. Can you provide a brief explanation of the specific locations marked by <regions> in this photo?
1. Whis is the object category of <regions>? Answer the question with a single word or phrase.
2. Could you tell me what is the object in <regions>? Answer the question with a single word or phrase.
3. What category best describes the area represented by <regions>? Answer the question with a single word or phrase.
4. Can you specify the type of object inside the region labeled by <regions>? Answer the question with a single word or phrase.
5. How would you label the area indicated by <regions> in the image? Answer the question with a single word or phrase.
6. Give a category label to the region outlined by <regions>. Answer the question with a single word or phrase.
7. Please identify the category of the object inside the <regions>. Answer the question with a single word or phrase.
8. Examine and determine the primary subject located within <regions>. Answer the question with a single word or phrase.
9. I need your help to assign an object category to the <regions>, please. Answer the question with a single word or phrase.
10. Evaluate the content of the region shown as <regions> and provide its category. Answer the question with a single word or phrase.